can you show me step by step how to solve this linear equation how do i graph it on a coordinate grid y=3x-4

Answer by JBarnum(2146)

plug in a number for x and solve for y
id say use these three for x:-1,0,1
y=3(-1)-4
y=-3-4
y=-7
y=3(0)-4
y=-4
y=3(1)-4
y=3-4
y=-1


so if x=-1,0,1 then y=-7,-4,-1
or stated in 3 points as
(-1,-7)(0,-4)(1,-1)


graphing is rather simple you can either use the points from above by plotting them and then drawing line or you can simply looke at the expression and know that y=mx+b where m=slope and b=y-intercept
y-intercept is at (0,-4)
and slope says for you to go up 3 and to the right 1 makinge second point (1,-1) and then draw line